# Hydrodynamics of a disk in a thin film of weakly nematic fluid subject to linear friction

### v1

@@ I Introduction
The diffusional motion of biological membrane components, including lipids and proteins, plays a crucial role in living systems, governing numerous biochemical functions. Early theoretical investigations into the Brownian motion of particles within a membrane date back to Saffman and Delbrück, who determined the hydrodynamic resistance experienced by a slender disk in motion within a thin sheet of viscous fluid representing a model of a lipid bilayer [1, 2]. In their seminal works, they supposed that the viscosity of the fluid representing the membrane is much larger than the viscosity of a surrounding exterior liquid. Later Hughes et al. [3] addressed the scenario involving arbitrary viscosities, providing exact expressions for resistance coefficients through techniques of dual integral equations. An extension of the Saffman theory was proposed beyond conventional hydrodynamic theories by introducing a generic model involving a protein coupled to a nonconserved scalar order parameter, such as chain stretching. [4]. More recently, the translational motion of a disk embedded in a nearly inviscid Langmuir film was revisited [5]. On the experimental side, progress in fluorescence imaging techniques enabled a monitoring of the dynamics and interactions of proteins within living cells [6, 7, 8, 9].
@@
The rheological properties of actual membranes typically deviate from Newtonian behavior. In various biologically relevant scenarios, different types of particles encounter the challenge of navigating through membranes characterized by complex properties [13]. Levine et al. extended the model to encompass viscoelastic films embedded within an infinitely extended viscous liquid, addressing both in- and out-of-plane dynamics [14, 15, 16]. Furthermore, hydrodynamic forces exerted on a laterally moving two-dimensional liquid domain within a supported fluid membrane, considering the concept of “odd viscosity”, were investigated in diverse contexts by Hosaka et al. [17, 18, 19]. More recently, the hydrodynamic aggregation of membrane-bound inclusions was examined under the influence of non-Newtonian surface rheology [20].
@@
In this study, we will utilize a two-dimensional Fourier transform method to examine how a disk behaves within a two-dimensional fluid film of uniaxial anisotropy. This technique, widely employed in studying elastic sheets and flexible membranes, capitalizes on the symmetry that exists across horizontal planes. The central idea involves setting boundary conditions for the velocity on the perimeter of the disk and there determining the force density as an unknown parameter. Previously, this approach has been successfully employed to analyze dynamics near various surfaces, including walls and interfaces [33, 34, 35, 36], as well as to understand the behavior of elastic membranes undergoing different modes of shear and bending deformation [37, 38, 39, 40, 41, 42, 43].
@@
The remainder of the presentation is organized as follows. In Sec. II, we provide an overview of the continuum description at low Reynolds numbers, employed to characterize the dynamics of the surrounding two-dimensional anisotropic sheet of ordered nematic fluid. In this context, we utilize Fourier transformations for the hydrodynamic fields. From the inverse Fourier transformation, we obtain in Sec. III perturbative analytical solutions for the flow field induced by a disk laterally translating in a weakly nematic fluid sheet. Next, in Sec. IV, we evaluate the impact of fluid anisotropy on the translational resistance coefficients of the disk laterally surrounded by the nematic fluid. Additionally, we present asymptotic expressions of the resistance coefficients for the limiting cases of weak and strong linear friction with a supporting substrate or other types of surroundings. In Sec. V, we compare our theoretical predictions with finite-element simulations. Following this comparison, a thorough analysis of the results is provided in Sec. VI. Conclusions are added in Sec. VII.
@@ II.1 Generalized Stokes hydrodynamics
Flows at small scales are characterized by low Reynolds numbers, where inertial forces are subordinate to viscous forces. Under these conditions, the dynamics of the surrounding fluid are predominantly governed by the Stokes equation. We here study an extension of this equation to take into account possible friction with the surrounding. In reality, such friction may arise from a substrate and possibly by an additional cover slide confining a thin, basically two-dimensional fluid layer, or, possibly and to some approximation, from the friction with a surrounding fluid. Consequently, the flow equation can be generally expressed as
$\nabla_{j}\sigma_{ij}(\bm{r})+M_{ij}v_{j}(\bm{r})=f_{i}(\bm{r})\,.$ (1)
[…]
For a simple isotropic fluid, $\bm{\widetilde{\sigma}}(\bm{r})$ introduces viscous dissipation into the dynamic equations, encompassing two viscosity parameters — one associated with shear viscosity and the other with dissipation during volume changes. Instead, we here address a uniaxially symmetric, anisotropic fluid medium. We start from the corresponding symmetry-based expressions for a conventional nematic liquid crystal deep within the nematic phase [46, 47]. The local orientation of the axis of uniaxial order is defined by the normalized director field $\bm{\hat{n}}(\bm{r})$.
Generally, the director field is dynamic, changing over time, for example, due to fluid flows or the action of external fields. Spatial variations in $\bm{\hat{n}}(\bm{r})$ require deformational energy and contribute to the stress tensor $\bm{\widetilde{\sigma}}(\bm{r})$.
In our considerations, we examine a scenario of the director being consistently and uniformly oriented along a single global axis [48, 49, 50]. This alignment can be achieved, for instance, through the influence of a strong aligning external electric field (under insulating conditions) or a magnetic field [51, 52, 53]. When $\bm{\hat{n}}(\bm{r})$ is perfectly aligned with the field, there are not any associated gradients contributing to the stress tensor.
Hence, the stress tensor $\bm{\widetilde{\sigma}}(\bm{r})$ is simplified to the pure dissipative stress resulting from gradients in the fluid flow [52, 53],
@@
following Ref. 54.
From now on, we do not explicitly mention the dependence of the fields on $\bm{r}$ any longer. Overall, we incorporate the common assumption made in similar analyses of fluid flows, namely, local volume conservation and constant density of the fluid. As a result, the continuity equation simplifies to
$\bm{\nabla}\cdot\bm{v}=0\,.$ (5)
@@
In the subsequent discussion, we assume that the friction tensor is symmetric. We can thus study the situation in its principal-axes-frame, where it adopts the form
@@
Inserting Eq. (6) into Eq. (1) and subtracting $\left(\nu_{3}+\nu_{5}\right)\bm{\nabla}\left(\bm{\nabla}\cdot\bm{v}\right)=0$ from the resulting equations, we obtain
@@ II.2 Fourier representation
To obtain the solution for the flow velocity and pressure fields, a two-dimensional Fourier transformation along both the $x$ and $y$ directions is employed. We define the forward Fourier transformation of a given function $g(\bm{\rho})$ as
$\widetilde{g}(\bm{k})=\mathscr{F}\left\{g(\bm{\rho})\right\}=\int_{\mathbb{R}^{2}}g(\bm{\rho})\,e^{-i\bm{k}\cdot\bm{\rho}}\,\mathrm{d}^{2}\bm{\rho}\,,$ (9)
@@
$g(\bm{\rho})=\mathscr{F}^{-1}\left\{\widetilde{g}(\bm{k})\right\}=\frac{1}{\left(2\pi\right)^{2}}\int_{\mathbb{R}^{2}}\widetilde{g}(\bm{k})\,e^{i\bm{k}\cdot\bm{\rho}}\,\mathrm{d}^{2}\bm{k}\,.$ (10)
In these expressions, $\bm{\rho}=(x,y)$ denotes the position vector. Furthermore, we introduce the wavenumber $k=|\bm{k}|$ and the unit wavevector $\bm{\hat{k}}=\bm{k}/k$. We employ a polar coordinate system to express $\bm{\hat{k}}=(\hat{k}_{x},\hat{k}_{y})=(\cos\phi,\sin\phi)$.
@@
Introducing the Green’s function $\widetilde{\bm{\mathcal{G}}}(k,\phi)$, the fluid velocity in Fourier space can be expressed as
$\widetilde{\bm{v}}(k,\phi)=\widetilde{\bm{\mathcal{G}}}(k,\phi)\cdot\widetilde{\bm{f}}(k,\phi)\,.$ (12)
@@ III Hydrodynamic flow fields
We consider the steady motion of a circular disk moving in a two-dimensional anisotropic sheet of a globally aligned uniaxial fluid, specifically, a nematic liquid crystal. The disk of radius $a$ at the considered time $t$ moves with speed $V$. No-slip boundary conditions for the flow field $\bm{(}\rho,\theta)$ of the fluid apply at the perimeter of the disk, They ensure $\bm{v}(\rho=a,\theta)=V_{\parallel}\bm{\hat{x}}+V_{\perp}\bm{\hat{y}}$, that is, the fluid acquires identical flow velocity over the complete perimeter of the disk. Without loss of generality, we here consider the time $t$ at which the center of the disk is located at the origin of our coordinate frame. Leveraging the linearity of the governing equations of motion, we can independently analyze the motion parallel and perpendicular to the nematic director $\bm{\hat{n}}\parallel\bm{\hat{x}}$.
@@
The coefficients $\bm{f}_{n}$ are to be determined based on the given boundary condition. The overall force exerted by the disk is later derived through integration of $\bm{f}(\rho,\theta)$ over the area of the fluid, that is, over the area of $\rho\geq a$. Evidently, it is found solely from the zeroth-order Fourier mode, $\langle\bm{F}\rangle=2\pi a\,\bm{f}_{0}$. Following a straightforward algebraic manipulation, the force density in Eq. (17) can be expressed in Fourier space as
@@ III.1 Velocity field
Our central step is to determine the force density distribution around the perimeter of the disk that there satisfies the boundary conditions for the velocity field. Since the governing equations of motion exhibit a linear nature, it is adequate to focus on the predominant two modes of motion, that is, parallel and perpendicular to the director. Here, we impose a no-slip boundary condition for the velocity field at the perimeter of the disk. Thus, for the velocity $\bm{V}=V\bm{\hat{V}}$ of the disk, the radial and azimuthal components of the velocity field at $\rho=a$ are given by $v_{\rho}=V\cos\theta$ and $v_{\theta}=-V\sin\theta$ for motion parallel to the director ($\bm{\hat{V}}\|\bm{\hat{x}}$). Similarly, they are given by $v_{\rho}=V\sin\theta$ and $v_{\theta}=V\cos\theta$ for motion perpendicular to the director ($\bm{\hat{V}}\|\bm{\hat{y}}$). Matching the Fourier modes at the surface of the disk shows that only even Fourier modes contribute to the force density.
By inserting Eq. (18) into Eq. (12) and applying inverse Fourier transformation, the velocity field can be expressed in real space as [56]
@@
Subsequently, by employing the Fourier representation of the force density as provided in Eq. (18), the velocity field can be expressed as an infinite double sum of the form,
@@
with the dimensionless numbers ${w}=a/\rho$ and $\sigma$ implying that
$B=\left(\frac{\rho}{\sigma}\right)^{2}\left(\alpha_{\parallel}^{2}\sin^{2}\phi+\alpha_{\perp}^{2}\cos^{2}\phi\right).$ (24)
Here, we introduced the change of variables $u=k\rho$. ${w}\leq 1$ marks the fluid region outside the disk. We recall that the index $m$ is employed in the context of the inverse Fourier transformation of the velocity field, whereas the index $n$ corresponds to the Fourier representation of the unknown force density.
In the following, we set for motion parallel to the director $\bm{f}_{2n}^{\parallel}=\left(c_{2n}^{\parallel},-id_{2n}^{\parallel}\right)^{\top}$, with $(\cdot)^{\top}$ denoting the transpose operator. It follows that $c_{-2n}^{\parallel}=c_{2n}^{\parallel}$ and $d_{-2n}^{\parallel}=-d_{2n}^{\parallel}$, implying that $d_{0}^{\parallel}=0$. For motion perpendicular to the director, we set $\bm{f}_{2n}^{\perp}=\left(-ic_{2n}^{\perp},d_{2n}^{\perp}\right)^{\top}$ so that $c_{-2n}^{\perp}=-c_{2n}^{\perp}$ and $d_{-2n}^{\perp}=d_{2n}^{\perp}$, implying that $c_{0}^{\perp}=0$. These coefficients are selected as real numbers. Their values are determined from the specified boundary conditions as detailed below.
@@
$\bm{v}(\rho,\theta)$ $=\frac{a}{\nu_{3}}\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}\left(U_{mn}\,\bm{\hat{x}}+V_{mn}\,\bm{\hat{y}}\right),$ (28)
wherein
$U_{mn}$ $=\phantom{+}\frac{2}{\pi}\int_{0}^{2\pi}\mathcal{M}_{mn}(\theta,\phi)\,\Phi_{mn}(\rho,\phi)\,\sin\phi\,\mathrm{d}\phi\,,$ (29a)
$V_{mn}$ $=-\frac{2}{\pi}\int_{0}^{2\pi}\mathcal{M}_{mn}(\theta,\phi)\,\Phi_{mn}(\rho,\phi)\,\cos\phi\,\mathrm{d}\phi\,,$ (29b)
where
$\mathcal{M}_{mn}(\theta,\phi)=\xi_{mn}\cos\left(2m(\phi-\theta)\right)Z_{2n}(\phi)\,,$ (30)
@@
As evident from Eqs. (23) and (III.1), determining the velocity field involves the assessment of a double integral. Due to the delicate nature of analytical evaluation, especially in the general case, we adopt the following approach. The integral over the scaled wavenumber, as defined by Eq. (23), allows for exact evaluation, yielding an analytical expression. However, when calculating the ensuing integrals over $\phi$ outlined in Eqs. (III.1), we resort to approximations.
@@
where $I_{2n}$ and $K_{2n}$ represent the modified Bessel functions (also known as the hyperbolic Bessel functions) of the first and second kind, respectively. Denoted as $\mathcal{P}\Psi_{mn}$, the principal part of the series expansion of $\Psi_{mn}$ around $\sigma=0$ refers to its component involving negative powers [57]. To clarify, the principal part corresponds to the segment of the Laurent series expansion of $\Psi_{mn}$ characterized by negative exponents. This result holds only when ${w}\leq 1$, which corresponds to the area outside the disk.

### v2

@@ I Introduction
The diffusional motion of biological membrane components, including lipids and proteins, plays a crucial role in living systems, governing numerous biochemical functions. Notable examples include signal transduction, where the movement of receptor proteins and their interactions with ligands facilitate the transmission of signals across the cell membrane [1, 2], and cellular communication, which relies on the diffusion of molecules like neurotransmitters and hormones across synaptic and cellular membranes [3, 4]. Early theoretical investigations into the Brownian motion of particles within a membrane date back to Saffman and Delbrück, who determined the hydrodynamic resistance experienced by a slender disk in motion within a thin sheet of viscous fluid representing a model of a lipid bilayer [5, 6] In their seminal works, they assumed that the viscosity of the fluid representing the membrane is much greater than that of the surrounding exterior liquid, defining the limit $\operatorname{Bq}\gg 1$ for the Boussinesq number $\operatorname{Bq}=\eta_{\mathrm{S}}/a\eta$. Here, $\eta_{\mathrm{S}}$ is the surface viscosity, $\eta$ is the viscosity of the exterior fluid, and $a$ is the radius of the disk. Later Hughes et al. [7] addressed the scenario involving arbitrary viscosities, providing exact expressions for resistance coefficients through techniques of dual integral equations. The problem with finite subphase depth was subsequently addressed by Stone and Ajdari [8], partially through numerical methods. For infinite depth and vanishing surface viscosity ($\operatorname{Bq}\to 0$), the translational drag coefficient was found to be 50% larger than the drag coefficient for a fluid with a free surface [8, 9]. More recently, the translational motion of a disk embedded in a nearly inviscid Langmuir film, marking the limit $\operatorname{Bq}\ll 1$, was revisited [10]. By utilizing the reciprocal theorem [11] in a fluid domain tailored to the asymptotic topology of the problem under investigation, the leading-order correction to the drag coefficient has been analytically obtained. An extension of the Saffman theory was proposed beyond conventional hydrodynamic theories by introducing a generic model involving a protein coupled to a nonconserved scalar order parameter, such as chain stretching. [12]. In that model, the diffusion coefficient is found to scale inversely with protein radius when the protein is much larger than the order parameter correlation length. This contrasts with the hydrodynamic Saffman-Delbrück theory, which predicts a logarithmic dependence of the diffusion coefficient on protein radius. The model is observed to align more closely with experimental findings that have measured protein diffusion coefficients, revealing a stronger correlation with protein radius [13, 14, 15]. On the experimental side, progress in fluorescence imaging techniques enabled a monitoring of the dynamics and interactions of proteins within living cells [16, 17, 18, 19].
@@
The rheological properties of actual membranes typically deviate from Newtonian behavior. In various biologically relevant scenarios, different types of particles bound to membranes often encounter the challenge of navigating through membranes characterized by complex properties [23]. Levine et al. extended the classical model for extended bodies embedded within infinitely extended viscous films and membranes, addressing both in- and out-of-plane dynamics [24, 25]. It has been found that for large inclusions, both rotation and motion perpendicular to the long axis demonstrate purely local drag behavior, where the drag coefficient is algebraic in terms of the dimensions of the elongated object. Moreover, hydrodynamic forces acting on laterally moving inclusions within a two-dimensional liquid domain confined by a supported fluid membrane, considering the concept of “odd viscosity”, have been explored in various contexts by Hosaka et al. [26, 27, 28]. The authors report that a disk of finite size moving in the two-dimensional fluid does not only experience a dissipative drag force that opposes its motion, but also a non-dissipative lift force. More recently, the hydrodynamic aggregation of membrane-bound inclusions was examined under the influence of non-Newtonian surface rheology [29]. In that case, a corrective force arises due to the interplay between membrane flow and non-constant viscosity. This suggests a mechanism for biologically favorable protein aggregation within membranes.
@@
The dynamics of colloidal particles immersed in a three-dimensional liquid crystal has been extensively studied over many years [42, 43, 44, 45]. Various types of anchoring and alignment of liquid crystal molecules on particle surfaces can elastically distort an otherwise spatially uniform alignment of the director field [46]. The axisymmetric flow field around a sphere dragged along the director of an undisturbed, aligned nematic liquid crystal has been theoretically studied, particularly when neglecting one viscosity coefficient, yielding analytical expressions for the frictional drag acting on the sphere in a quiescent liquid crystal [47]. Later, a general solution for arbitrary orientation and viscosity coefficients was obtained numerically [48]. Additionally, closed-form analytical formulas derived from conservation laws for nematic liquid crystals have been developed using techniques of Fourier transformation [49, 50]. Alongside, the Stokes drag of a spherical particle in various nematic environments has been investigated through computer simulations [51, 52, 53, 54]. Experimental studies have shown that self-diffusion in a nematic liquid crystal follows a generalized Stokes-Einstein relation, with the effective diffusion coefficient along the oriented far-field director typically greater than perpendicular to it under the given circumstances [55, 56, 57, 58].
@@
In this study, we utilize a method of two-dimensional Fourier transformation to examine how a disk behaves within a two-dimensional fluid film of uniaxial anisotropy. Previously, this method has been successfully employed to analyze dynamics near various surfaces, including walls and interfaces [59, 60, 61, 62, 63], as well as to understand the behavior of elastic membranes undergoing different modes of shear and bending deformation [64, 65, 66, 67, 68, 69, 70, 71]. Thus, widely employed in studying elastic sheets and flexible membranes, this technique capitalizes on the symmetry that exists across horizontal planes. The central idea involves setting boundary conditions for the velocity on the perimeter of the disk and there determining the force density as an unknown parameter. There are well-established approaches available in classical textbooks on fluid mechanics that provide analogous methods [72, 73, 74]. However, they are primarily applicable to isotropic fluids.
@@
The remainder of the presentation is organized as follows. In Sec. II, we provide an overview of the continuum description at low Reynolds numbers, employed to characterize the dynamics of the surrounding two-dimensional anisotropic sheet of ordered nematic fluid. In this context, we utilize Fourier transformations for the hydrodynamic fields. From the inverse Fourier transformation, we obtain in Sec. III perturbative analytical solutions for the flow field induced by a disk laterally translating in a weakly nematic fluid sheet. Next, in Sec. IV, we evaluate the impact of fluid anisotropy on the translational resistance coefficients of the disk laterally surrounded by the nematic fluid. Additionally, we present asymptotic expressions of the resistance coefficients for the limiting cases of weak and strong linear friction with a supporting substrate or other types of surroundings. In Sec. V, we compare our theoretical predictions with finite-element simulations. Following this comparison, an analysis of the results is provided in Sec. VI. Conclusions are added in Sec. VII. Besides, Appendix A contains a summary of the key parameters introduced in this manuscript for convenient reference. Appendix B presents details of calculating a non-trivial infinite integral in Sec. III.
@@ II.1 Generalized Stokes hydrodynamics
For flows at low Reynolds numbers, viscous contributions dominate inertial ones. Under these conditions, the dynamics of the surrounding fluid are predominantly governed by the Stokes equation. We here study an extension of this equation to take into account possible friction with the surrounding. In reality, such friction may arise from a substrate and possibly by an additional cover slide confining a thin, basically two-dimensional fluid layer, or, possibly and to some approximation, from the friction with a surrounding fluid. Consequently, the flow equation can be generally expressed as
$\nabla_{j}\sigma_{ij}(\boldsymbol{r})+M_{ij}v_{j}(\boldsymbol{r})=f_{i}(\bm{r})\,.$ (1)
In this equation, $\sigma_{ij}(\boldsymbol{r})=p(\boldsymbol{r})\,\delta_{ij}+\widetilde{\sigma}_{ij}(\boldsymbol{r})$ represents the components of the stress tensor, where $p(\boldsymbol{r})$ is the pressure field, $\delta_{ij}$ is the Kronecker delta, $f_{i}(\bm{r})$ in two dimensions signifies a force density acting on the fluid, and summation over repeated indices is implied following Einstein’s notation. $M_{ij}$ represent the coefficients of linear friction with the substrate or top and bottom surroundings. $\boldsymbol{v}(\boldsymbol{r})$ is the velocity field. (In a different context, $\bm{M}$ played the role of a permeability tensor of SI unit kg/s m2, particularly when considering flows through porous media [75].)
For a simple isotropic fluid, $\bm{\widetilde{\sigma}}(\boldsymbol{r})$ introduces viscous dissipation into the dynamic equations, encompassing two viscosity parameters — one associated with shear viscosity and the other with dissipation during volume changes. Instead, we here address a uniaxially symmetric, anisotropic fluid medium. We start from the corresponding symmetry-based expressions for a conventional nematic liquid crystal deep within the nematic phase [76, 77]. The local orientation of the axis of uniaxial order is defined by the normalized director field $\bm{\hat{n}}(\bm{r})$.
Generally, the director field is dynamic, changing over time, for example, due to fluid flows or the action of external fields. Spatial variations in $\bm{\hat{n}}(\bm{r})$ require deformational energy and contribute to the stress tensor $\bm{\widetilde{\sigma}}(\boldsymbol{r})$.
In our considerations, we examine a scenario of the director being consistently and uniformly oriented along a single global axis [78, 48, 79]. This alignment can be achieved, for instance, through the influence of a strong aligning external electric field (under insulating conditions) or a magnetic field [80, 81, 82]. When $\bm{\hat{n}}(\boldsymbol{r})$ is perfectly aligned with the field, there are not any associated gradients contributing to the stress tensor.
Hence, the stress tensor $\bm{\widetilde{\sigma}}(\boldsymbol{r})$ is simplified to the pure dissipative stress resulting from gradients in the fluid flow [81, 82],
@@
following Ref. 83.
From now on, we do not explicitly mention the dependence of the fields on $\boldsymbol{r}$ any longer. Overall, we incorporate the common assumption made in similar analyses of fluid flows, namely, local volume conservation and constant density of the fluid. As a result, the continuity equation simplifies to
$\boldsymbol{\nabla}\cdot\boldsymbol{v}=0\,.$ (5)
@@
In the subsequent discussion, we examine the scenario of one of the principal axes of the friction tensor aligning with the nematic director. Thus, the friction tensor in our representation adopts a diagonal form
@@
Inserting Eq. (6) into Eq. (1) and subtracting $\left(\nu_{3}+\nu_{5}\right)\boldsymbol{\nabla}\left(\boldsymbol{\nabla}\cdot\boldsymbol{v}\right)=0$ from the resulting equations, we obtain
@@ II.2 Fourier representation
To obtain the solution for the flow velocity and pressure fields, a two-dimensional Fourier transformation along both the $x$ and $y$ directions is employed. We define the forward Fourier transformation of a given function $g(\boldsymbol{\rho})$ as
$\widetilde{g}(\bm{k})=\mathscr{F}\left\{g(\boldsymbol{\rho})\right\}=\int_{\mathbb{R}^{2}}g(\boldsymbol{\rho})\,e^{-i\bm{k}\cdot\boldsymbol{\rho}}\,\mathrm{d}^{2}\boldsymbol{\rho}\,,$ (9)
@@
$g(\boldsymbol{\rho})=\mathscr{F}^{-1}\left\{\widetilde{g}(\bm{k})\right\}=\frac{1}{\left(2\pi\right)^{2}}\int_{\mathbb{R}^{2}}\widetilde{g}(\bm{k})\,e^{i\bm{k}\cdot\boldsymbol{\rho}}\,\mathrm{d}^{2}\bm{k}\,.$ (10)
In these expressions, $\boldsymbol{\rho}=(x,y)$ denotes the position vector. Furthermore, we introduce the wavenumber $k=|\bm{k}|$ and the unit wavevector $\bm{\hat{k}}=\bm{k}/k$. We employ a polar coordinate system to express $\bm{\hat{k}}=(\hat{k}_{x},\hat{k}_{y})=(\cos\phi,\sin\phi)$.
@@
For the common representatives of liquid crystals MBBA (N-(4-methoxybenzylidene)-4-butylaniline) and 5CB (4-cyano-4’-pentylbiphenyl), we estimate a typical value of $A\approx 4$ based on literature data [84, 83]. Physically, it is important to highlight that A must meet the condition $A>-4$, which is crucial for thermodynamic stability. This condition is derived from the necessity that $A>\nu_{2}/\nu_{3}-4$. Given that both $\nu_{2}$ and $\nu_{3}$ must be positive, it follows that $A>-4$. Below, we demonstrate how this parameter serves as a perturbation parameter in obtaining approximate solutions for the hydrodynamic flow fields.
The solution for the velocity field in Fourier space can be obtained by taking the Fourier transformations of Eqs. (8). Subsequently, we apply the divergence-free condition of the velocity field to eliminate the pressure field from these equations. Introducing the Green’s function $\widetilde{\bm{\mathcal{G}}}(k,\phi)$, the fluid velocity in Fourier space can be expressed as
$\widetilde{\boldsymbol{v}}(k,\phi)=\widetilde{\bm{\mathcal{G}}}(k,\phi)\cdot\widetilde{\bm{f}}(k,\phi)\,.$ (12)
@@ III Hydrodynamic flow fields
Figure 1: Schematic illustration of the quantities involved for the two-dimensional disk under consideration. The circular disk of radius $a$ moves in a two-dimensional anisotropic sheet of a globally aligned uniaxial fluid. The axis of anisotropy, characterized by the nematic director $\hat{\bm{n}}$, is aligned permanently along the $x$-direction. Our disk moves with velocity components $V_{\parallel}$ parallel and $V_{\perp}$ perpendicular to the director. We work within the frame of reference anchored to the disk, so that the origin of the coordinate frame remains at the center of the disk. The frictional effect with a supporting substrate or top and bottom surroundings is represented by a linear friction tensor.
We consider the steady motion of a circular disk moving in a two-dimensional anisotropic sheet of a globally aligned uniaxial fluid, specifically, a nematic liquid crystal. The disk of radius $a$ moves with speed $V$. No-slip boundary conditions for the flow field $\boldsymbol{(}\rho,\theta)$ of the fluid apply at the perimeter of the disk. They ensure $\boldsymbol{v}(\rho=a,\theta)=V_{\parallel}\bm{\hat{x}}+V_{\perp}\bm{\hat{y}}$, that is, the fluid acquires identical flow velocity over the complete perimeter of the disk. Our approach of solution concentrates on the comoving frame of reference that permanently anchors its origin to the center of the disk. Leveraging the linearity of the governing equations of motion, we can independently analyze the motion parallel and perpendicular to the nematic director $\bm{\hat{n}}\parallel\bm{\hat{x}}$. Figure 1 provides an illustration of the quantities involved in the geometry. It is important to highlight that our attention is primarily placed on the dynamics of the considered two-dimensional fluid. We do not describe the response of a possible environment surrounding this thin fluid film, for instance, a supporting substrate or a fluid or gaseous component beneath or above the fluid film and the disk. The influence of the surrounding environment, the dynamics of which we do not include in our consideration, is represented by the friction tensor specified by Eq. (7).
@@
The coefficients $\bm{f}_{n}$ are to be determined based on the given boundary condition. We note that the dimension of the coefficients $\bm{f}_{n}$ is not identical to those of the force density $\bm{f}$, because the dimension of the delta function is of inverse length. The overall force exerted by the disk is derived later through integration of $\bm{f}(\rho,\theta)$ over the area of the fluid, that is, over the area of $\rho\geq a$. Evidently, it is found solely from the zeroth-order Fourier mode, $\langle\bm{F}\rangle=2\pi a\,\bm{f}_{0}$. Following a straightforward algebraic manipulation, the force density in Eq. (17) can be expressed in Fourier space as
@@ III.1 Velocity field
Our central step is to determine the force density distribution around the perimeter of the disk that there satisfies the boundary conditions for the velocity field. Since the governing equations of motion exhibit a linear nature, it is adequate to focus on the predominant two modes of motion, that is, parallel and perpendicular to the director. Here, we impose a no-slip boundary condition for the velocity field at the perimeter of the disk. Thus, for the velocity $\bm{V}=V\bm{\hat{V}}$ of the disk, the components of the velocity field of the fluid at $\rho=a$ are given by $v_{x}(\rho=a)=V$ and $v_{y}(\rho=a)=0$ for motion parallel to the director ($\bm{\hat{V}}\|\bm{\hat{x}}$). Similarly, they are given by $v_{x}(\rho=a)=0$ and $v_{y}(\rho=a)=V$ for motion perpendicular to the director ($\bm{\hat{V}}\|\bm{\hat{y}}$). By matching the Fourier modes at the surface of the disk, we obtain the solution for the force density in terms of a Fourier series.
It is important to clarify that we project vector quantities onto Cartesian coordinates, in accordance with the geometry of the system. Conversely, we express positions both in real and Fourier spaces using polar coordinates. By applying inverse Fourier transformation, the velocity field can be expressed in real space as [86]
@@
Since the boundary conditions prescribed at the surface of the disk correspond to constant velocities, the velocity field given by Eq. (19), when evaluated at $\rho=a$, is
[…]
for $m\geq 0$. Accordingly, only the zeroth mode matches the velocity of the disk at $\rho=a$, while all other modes are set to zero.
Subsequently, utilizing Eqs. (12) and (18), we find that the $\phi$-integration in Eq. (20) vanishes, if $m-n$ is an odd number. Since Eqs. (21) represent a system of linear equations for the unknown values of $\bm{f}_{n}$, it follows that all odd Fourier modes in the expression of the force density must vanish. Consequently, only the even modes will persist in the series expansion of the velocity field as well.
Then, the velocity field can be expressed as an infinite double sum of the form,
@@
with the dimensionless numbers ${w}=a/\rho$ and
$\sigma=\frac{\rho}{B^{1/2}}\left(\alpha_{\parallel}^{2}\sin^{2}\phi+\alpha_{\perp}^{2}\cos^{2}\phi\right)^{1/2}.$ (25)
We note that $B$ is a function of $\phi$, as previously defined below Eq. (14). Moreover, we introduced the change of variables $u=k\rho$. ${w}\leq 1$ marks the fluid region outside the disk. We recall that the index $m$ is employed in the context of the inverse Fourier transformation of the velocity field, whereas the index $n$ corresponds to the Fourier representation of the unknown force density.
In the following, we set for motion parallel to the director $\bm{f}_{2n}^{\parallel}=\left(c_{2n}^{\parallel},-id_{2n}^{\parallel}\right)^{\top}$, with $(\cdot)^{\top}$ denoting the transpose operator. It follows that $c_{-2n}^{\parallel}=c_{2n}^{\parallel}$ and $d_{-2n}^{\parallel}=-d_{2n}^{\parallel}$, implying that $d_{0}^{\parallel}=0$. For motion perpendicular to the director, we set $\bm{f}_{2n}^{\perp}=\left(-ic_{2n}^{\perp},d_{2n}^{\perp}\right)^{\top}$ so that $c_{-2n}^{\perp}=-c_{2n}^{\perp}$ and $d_{-2n}^{\perp}=d_{2n}^{\perp}$, implying that $c_{0}^{\perp}=0$. These coefficients are selected as real numbers. Their values are determined from the specified boundary conditions, as detailed below.
@@
$\boldsymbol{v}{{}^{k}}(\rho,\theta)$ $=\frac{a}{\nu_{3}}\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}\left({U_{mn}^{k}\,\bm{\hat{x}}+V_{mn}^{k}\,\bm{\hat{y}}}\right),$ (29)
for $k\in\{\parallel,\perp\}$, wherein
${U_{mn}^{k}}$ $=\phantom{+}\frac{2}{\pi}\int_{0}^{2\pi}{\mathcal{M}_{mn}^{k}}(\theta,\phi)\,\Phi_{mn}(\rho,\phi)\,\sin\phi\,\mathrm{d}\phi\,,$ (30a)
${V_{mn}^{k}}$ $=-\frac{2}{\pi}\int_{0}^{2\pi}{\mathcal{M}_{mn}^{k}}(\theta,\phi)\,\Phi_{mn}(\rho,\phi)\,\cos\phi\,\mathrm{d}\phi\,,$ (30b)
@@
As evident from Eqs. (24) and (30), determining the velocity field involves the assessment of a double integral. Due to the delicate nature of analytical evaluation, especially in the general case, we adopt the following approach. The integral over the scaled wavenumber, as defined by Eq. (24), allows for exact evaluation, yielding an analytical expression. However, when calculating the ensuing integrals over $\phi$ outlined in Eqs. (30), we resort to approximations.
@@
where $I_{2n}$ and $K_{2n}$ represent the modified Bessel functions (also known as the hyperbolic Bessel functions) of the first and second kind, respectively. Denoted as $\mathcal{P}\Psi_{mn}$, the principal part of the series expansion of $\Psi_{mn}$ around $\sigma=0$ refers to its component involving negative powers [87]. To clarify, the principal part corresponds to the segment of the Laurent series expansion of $\Psi_{mn}$ characterized by negative exponents. This result only holds for ${w}\leq 1$, which corresponds to the area outside the disk. A detailed analytical solution is provided in Appendix B.